The line is perpendicular to 3x - 2y =  10 and passes
through the point (1, 1).

Answer by ewatrrr(24785)


Hi


3x - 2y = 10 OR y = (1/2)x - 5 |m = -2 is slope of line ⊥ to it
New Line:
y = -2x + b |using order pair Pt(1,1) to solve for b
1 = -2 + b
3 = b
y = -2x + 3